OpenStudy (rayman5766):

What value of z makes the statement true? –(z + (–17)) = 12 + 17 A. –46 B. –12 C. 12 D. 46

OpenStudy (anonymous):

\[-(z+(-17))=12+17\]Plug in any of the answer choices for z. \[-(-12+(-17))=12+17\] Distribute the negative. \[12+17=12+17\] Which equals: \[29=29 \] So the answers is -12. Just plug each number in for z and see if it is equal on both sides.